Kirsten has 64 coins in her piggy bank.She has $9.25 in dimes and quarters. How many dimes and how many quarters does she have?
I am doing Singapore math. I am in 5th grade.
Please help!!

Answer by oberobic(2304)

With coin problems, you have to keep track of the number of coins and their values.
d = number of dimes
10d = value of the dimes in cents
q = number of quarters
25q = value of the quarters in cents
.
10d + 25q = 925 cents (to keep all the amounts in the same units of measure)
.
d + q = 64
so
d = 64 - q
.
substitute
.
10(64-q) + 25q = 925
640 -10q + 25q = 925
15q = 285
q = 19
.
d = 64-19 = 45
.
Check the value to be sure this is the answer.
10d = 10(45) = 450
25q = 25(19) = 475
450 + 475 = 925
Correct.
.
Answer: Kirsten as 19 quarters and 45 dimes.
.
Done.